OpenStudy (anonymous):

Simplify the given expression. square root of negative 6 end root times square root of negative 18

OpenStudy (espex):

Because they are both roots, and they are being multiplied together, you can put them under the same radical. Once there you can break apart the -18 into its components and you want to put the '-' with the 6 so that you can remove it from under the radical. \[\sqrt{-6} \times \sqrt{-18} \rightarrow \sqrt{-6 * -18} \rightarrow \sqrt{-6*-6*3} \rightarrow -6\sqrt{3}\]

OpenStudy (mathstudent55):

That's not how you handle roots of negative numbers. Every root of a negative number has to be changed to i.